# A highly efficient single-loop smoothing damped Newton method for large-scale bilevel hyperparameter optimization of SVC

### v1

@@ 5 Damped Newton method for (NLPEC)
2:If $\|F_{\epsilon}(r^{k})\|_{2}=0$, stop. Otherwise, go to S3.
3:Find a solution $d^{k}$ of the linear system by BICG-STAB [26]
$F_{\epsilon}(r^{k})+\Gamma^{k}d^{k}=0,$ (19)
@@
4:Let $s_{k}=\rho^{i_{k}}$, where $i_{k}$ is the smallest integer in {0,1,2,…} such that the following holds
$g_{\epsilon}(r^{k}+\rho^{i_{k}}d^{k})\leq g_{\epsilon}(r^{k})+\sigma\rho^{i_{k}}\nabla g_{\epsilon}(r^{k})^{\top}d^{k}.$
5:Let $r^{k+1}=r^{k}+s_{k}d^{k}$, and $k:=k+1$. Go to S2.
Here we would like to highlight that the algorithm (SDNM) is a single-loop algorithm, which fills into the traditional line-search update: $r^{k+1}=r^{k}+s_{k}d^{k}$, and $d^{k}$ is the search algorithm obtained by solving a linear equation (19) in an inexact way. Therefore, there is no subproblem that need to be further solved by optimization package. In other words, our proposed method SDNM is package independent.
@@
$J_{r}F_{\epsilon}(\bar{r})$ $=\left[\begin{array}[]{cc}\nabla^{2}_{vv}L_{\epsilon}(\bar{v},\bar{\lambda})&\nabla^{2}_{v\lambda}L_{\epsilon}(\bar{v},\bar{\lambda})\\ \nabla^{2}_{v\lambda}L_{\epsilon}(\bar{v},\bar{\lambda})^{\top}&\mathbf{0}\end{array}\right]$ (20)
$=\left[\begin{array}[]{cc}\nabla^{2}_{vv}L_{\epsilon}(\bar{v},\bar{\lambda})&-J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))^{\top}\\ -J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))&\mathbf{0}\end{array}\right]\in\mathbb{R}^{(2m+1)\times(2m+1)}.$
@@
By (21b), we can see that $d^{v}\in\mathcal{C}(\bar{v},\bar{\lambda})$. Together with Assumption 4.5, (22) implies that $d^{v}=0$. (21a) hence reduces to $J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))^{\top}d^{\lambda}=0.$ By the full-row rankness of $J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))$, $d^{\lambda}=0$ is the unique solution. Therefore, $J_{r}F_{\epsilon}(\bar{r})$ is nonsingular. The proof is finished. ∎
@@
Let ${r^{k}}$ be generated by Algorithm 1. Suppose Assumption 4.5 holds at the accumulation point $\bar{r}$ of ${r^{k}}$, then
  • (a)
    We have $\nabla g_{\epsilon}(\bar{r})=0$.
  • (b)
    If the accumulation point $\bar{r}$ satisfies $F_{\epsilon}(\bar{r})=0$, then ${r^{k}}$ converge to $\bar{r}$ Q-quadratically.
@@
So far, we have devised an efficient method called Smoothing Damped Newton method (SDNM) to solve the bilevel optimization problem with cross-validation (1) by first smoothing the derived (MPEC) to (NLPEC), and then using the damped Newton method to solve the smoothing problem (NLPEC) by its KKT conditions (14). Theorem 5.2 implies that, under certain assumption (Assumption 4.5), the proposed smoothing damped Newton method (SDNM) enjoys a quadratic convergence rate.
@@ 6 Numerical Experiments
1:Given $T$, split the data set into a subset $\Omega$ with $l_{1}$ points and a hold-out test set $\Theta$ with $l_{2}$ points. The set $\Omega$ is equally partitioned into $T$ pairwise disjoint subsets, one for each fold.
@@
3:Post-processing procedure. The regularization hyperparameter $\hat{C}$ is rescaled by a factor $T/(T-1)$. This gives the final classifier $\hat{w}$.
@@ 6.1 The performance of SDNM
In this subsection, we discuss about SDNM from the following three aspects: (i) Verification of Assumption 4.5; (ii) The influence of smoothing parameter $\epsilon$; (iii) The decrease of $\|F_{\epsilon}(r^{k})\|_{2}$.
@@
One can observe from Table 2 that for most data set, the outer loop reaches the stopping criteria within 30 iterations by SDNM. We define $A_{2}:=\bar{r}_{v}^{\top}\nabla^{2}_{vv}L_{\epsilon}(v,\lambda)\bar{r}_{v}$, and it can be seen that $A_{2}$ is always positive, which satisfies Assumption 4.5, implying that SDNM provides strict local minimizers of (NLPEC).
@@
No. Dataset $C$ $A_{2}$ $(k,iter)$
1 diabetes 3.65 990.78 $(\mathbf{12},812)$
2 heart 0.42 18.39 $(\mathbf{7},1058)$
3 australian 0.80 192.43 $(\mathbf{11},1780)$
4 german.number 15.01 867.54 $(\mathbf{14},576)$
5 ionosphere 6.37 733.99 $(\mathbf{7},466)$
6 sonar 4.61 411.16 $(\mathbf{8},360)$
7 phishing 3.56 2223.30 $(\mathbf{11},631)$
8 w1a 0.05 39.18 $(\mathbf{8},541)$
9 w2a 2.84 1141.97 $(\mathbf{5},350)$
10 w3a 0.29 715.60 $(\mathbf{9},1131)$
11 w4a 2.33 1655.50 $(\mathbf{6},701)$
12 w5a 0.10 78.92 $(\mathbf{11},1136)$
6.1.2 Smoothing parameter $\epsilon$
@@
Figure 1: $E_{tr}(\%)$ of each smoothing parameter $\epsilon$
Figure 2: $k$ (outer loop) of different smoothing parameter $\epsilon$’s
Generally speaking, when $\epsilon$ is small ($10^{-5}$ or $10^{-4}$), the results are better. When $\epsilon$ increases to $10^{-2}$, the number of iterations $(k,iter)$ increases significantly, leading to more time consumption. Moreover, it’s difficult to find a universal optimal value of $\epsilon$ for all the given datasets, and the solution of SDNM is a bit sensitive to the perturbations of $\epsilon$. Overall, the value of $\epsilon$ is $10^{-4}$ in our following test.
6.1.3 The decrease of $\|F_{\epsilon}(r^{k})\|_{2}$
Figure 3 demonstrates the typical decrease of $\|F_{\epsilon}(r^{k})\|_{2}$ for some data sets, where one can indeed observe the quadratic convergence rate of SDNM.
Figure 3: $\log_{10}\|F_{\epsilon}(r^{k})\|_{2}$ along iterations by SDNM
6.2 Comparison with other methods
We set the initial values of the parameters of SDNM as $\ \epsilon=10^{-4},\ C=1,\ \rho=0.5,\ \sigma=1\times{10}^{-4}$. We compare our method with the following two methods:
  • fmincon, in which we solve (MPEC) directly with a Matlab built-in function fmincon.
  • SGRM, in which we solve (MPEC) by Scholtes global relaxation method proposed in [18]. SNOPT [12] is used to solve the subproblem in each iteration $k$ (given $t_{k}>0$):
    $\min\limits_{v\in\mathbb{R}^{m+1}}$ $\quad F(v)$ (MPECr)
    $\mathrm{s.t.}$ $\quad G_{i}(v)\geq 0,\ \forall i=1,\cdots,m,$
    $\quad H_{i}(v)\geq 0,\ \forall i=1,\cdots,m,$
    $\quad G_{i}(v)H_{i}(v)\leq t_{k},\ \forall i=1,\cdots,m.$
@@ 6.2 Comparison with other methods
The three methods are summarized below in Figure 4.
Figure 4: The roadmap of the three methods
@@
    CPU time ($t$), which is the time of hyperparameter optimization solving by aforementioned methods with the given data set in Table 1.
@@
    Test error ($E_{tr}\left(\%\right)$) as defined by
    $E_{tr}=\frac{1}{l_{2}}\sum\limits_{(x_{i},y_{i})\in\Theta}\lvert{\rm sign}\left(x_{i}^{\top}w^{*}\right)-y_{i}\rvert$,
    where $\Theta$ is the test set.
@@
The results are reported in Table 3, where we mark the winners of $t$, $E_{tr}\left(\%\right)$ and $E_{CV}\left(\%\right)$ in bold. As shown in Table 3, SDNM performs better than other methods in terms of test error ($E_{tr}(\%)$) and CV error ($E_{CV}\left(\%\right)$) on almost all datasets, especially for large-scale datasets (No.8 to No.12 where $n$ = 300), implying that this method has superior generalization performance on classification prediction and solving the over-fitting problem.
@@
No. Dataset Method $(\beta_{1},\beta_{2})$ $C$ $t$ $t_{ratio}$ $E_{tr}$ $E_{CV}$
1 diabetes fmincon $(5400,1801)$ 9.80 16.5 2.3 23.5 $\mathbf{22.7}$
SGRM $(5400,1801)$ 1.50 8.7 1.2 23.7 23.7
SDNM $(1800,1801)$ 3.65 $\mathbf{7.3}$ 1.0 $\mathbf{23.3}$ 23.3
2 heart fmincon $(2700,901)$ 11.81 3.4 3.3 18.3 $\mathbf{16.0}$
SGRM $(2700,901)$ 1.62 1.7 1.6 17.5 18.0
SDNM $(900,901)$ 0.42 $\mathbf{1.1}$ 1.0 $\mathbf{15.8}$ 18.0
3 australian fmincon $(5400,1801)$ 4.83 17.7 1.7 14.6 $\mathbf{13.3}$
SGRM $(5400,1801)$ 1.46 15.9 1.5 14.6 14.0
SDNM $(1800,1801)$ 0.80 $\mathbf{10.4}$ 1.0 $\mathbf{13.3}$ 14.3
4 german. fmincon $(5400,1801)$ 25.15 18.3 2.9 25.3 $\mathbf{26.7}$
number SGRM $(5400,1801)$ 1.40 12.1 1.9 $\mathbf{24.7}$ 28.7
SDNM $(1800,1801)$ 15.01 $\mathbf{6.2}$ 1.0 25.3 $\mathbf{26.7}$
5 ionosphere fmincon $(4320,1441)$ 8.76 11.6 4.5 $\mathbf{2.7}$ 26.7
SGRM $(4320,1441)$ 1.50 5.5 2.1 4.5 $\mathbf{25.4}$
SDNM $(1440,1441)$ 6.37 $\mathbf{2.6}$ 1.0 $\mathbf{2.7}$ 26.7
6 sonar fmincon $(2700,901)$ 13.39 4.4 6.0 $\mathbf{27.6}$ $\mathbf{30.0}$
SGRM $(2700,901)$ 2.64 2.8 3.8 $\mathbf{27.6}$ $\mathbf{30.0}$
SDNM $(900,901)$ 4.61 $\mathbf{0.7}$ 1.0 $\mathbf{27.6}$ 30.7
7 phishing fmincon $(5400,1801)$ 5.64 18.0 3.4 6.4 $\mathbf{7.0}$
SGRM $(5400,1801)$ 1.54 17.4 3.3 6.4 8.3
SDNM $(1800,1801)$ 3.56 $\mathbf{5.3}$ 1.0 $\mathbf{6.2}$ $\mathbf{7.0}$
8 w1a fmincon $(8100,2701)$ 3.46 35.7 3.7 3.5 3.6
SGRM $(8100,2701)$ 1.50 231.6 24.0 3.3 3.8
SDNM $(2700,2701)$ 0.05 $\mathbf{9.7}$ 1.0 $\mathbf{2.5}$ $\mathbf{3.1}$
9 w2a fmincon $(8100,2701)$ 3.46 36.1 4.3 $\mathbf{3.2}$ 3.6
SGRM $(8100,2701)$ 1.51 137.1 16.5 $\mathbf{3.2}$ 3.6
SDNM $(2700,2701)$ 2.84 $\mathbf{8.3}$ 1.0 $\mathbf{3.2}$ $\mathbf{3.3}$
10 w3a fmincon $(8100,2701)$ 4.06 37.1 2.3 2.4 3.3
SGRM $(8100,2701)$ 1.50 250.8 15.5 2.4 3.3
SDNM $(2700,2701)$ 0.29 $\mathbf{16.2}$ 1.0 $\mathbf{2.0}$ $\mathbf{2.9}$
11 w4a fmincon $(8100,2701)$ 3.64 36.6 3.6 $\mathbf{3.1}$ $\mathbf{3.1}$
SGRM $(8100,2701)$ 1.48 236.0 22.9 $\mathbf{3.1}$ $\mathbf{3.1}$
SDNM $(2700,2701)$ 2.33 $\mathbf{10.3}$ 1.0 $\mathbf{3.1}$ $\mathbf{3.1}$
12 w5a fmincon $(8100,2701)$ 3.66 37.0 2.1 4.1 4.2
SGRM $(8100,2701)$ 1.52 257.2 14.9 4.1 4.4
SDNM $(2700,2701)$ 0.10 $\mathbf{17.3}$ 1.0 $\mathbf{2.1}$ $\mathbf{3.8}$
The scale of the optimization model that each method solve is also listed in the table, denoted as $(\beta_{1},\beta_{2})$, where $\beta_{1}$ represents the number of constraints and $\beta_{2}$ represents the number of dimensions of variable in the optimization problem of each methods, as shown in Table 4. It can be seen that the smoothing process of SDNM reduces the scale of the problem effectively while solving large-scale problems.
Table 4: Scale of three methods
Method Problem $(\beta_{1},\beta_{2})$
fmincon (MPEC) $(3m,m+1)$
SGRM (MPECr) $(3m,m+1)$
SDNM (NLPEC) $(m,m+1)$
In terms of the CPU time ($t$) summarized in Figure 5, SDNM takes the shortest time among the three methods in all data sets. In order to demonstrate how much the SDNM accelerates the optimization process, in Table 3, we show the time ratio ($t_{ratio}$), which shows that SNDM is about twice as fast as the better one of the existing approaches (fmincon and SGRM) generally. In particular, for large-scale datasets (No.8 to No.12 where $n$ = 300), the CPU time of SDNM is much less than that of other remaining methods. For example, dataset w4a (No.11) with sample data 450 and feature number 300, it can return a solution of hyperparameter in 10 seconds, which is over 20 times faster than SGRM and over 3 times faster than fmincon.
Figure 5: CPU time (t) of three methods

### v3

@@ 5.3 Damped Newton method for (NLPϵ)
2:while $\|F_{\epsilon}(r^{k})\|_{2}\neq 0$, do
3: Find a solution $d^{k}$ of the linear system by BICG-STAB
$\quad F_{\epsilon}(r^{k})+\Gamma^{k}d^{k}=0,$ (25)
@@
4: Let $s_{k}=\rho^{i_{k}}$, where $i_{k}$ is the smallest non-negative integer satisfying
$\quad g_{\epsilon}(r^{k}+\rho^{i_{k}}d^{k})\leq g_{\epsilon}(r^{k})+\sigma\rho^{i_{k}}\nabla g_{\epsilon}(r^{k})^{\top}d^{k},$
 where $g_{\epsilon}(r)=\frac{1}{2}\|F_{\epsilon}(r)\|_{2}^{2}$.
5: Let $r^{k+1}=r^{k}+s_{k}d^{k}$, and $k:=k+1$.
6:end while
7:Return the final iterate $r_{\rm opt}=r^{k}$.
@@
$J_{r}F_{\epsilon}(\bar{r})=\left[\begin{array}[]{cc}\nabla^{2}_{vv}L_{\epsilon}(\bar{v},\bar{\lambda})&-J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))^{\top}\\ -J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))&\mathbf{0}\end{array}\right].$ (26)
@@
By (27b), we can see that $d^{v}\in\mathcal{C}(\bar{v},\bar{\lambda})$. Together with Assumption 5.5, (28) implies that $d^{v}=0$. (27a) hence reduces to $J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))^{\top}d^{\lambda}=0.$ By the full row rank of $J_{v}\Phi_{\epsilon}(G(\bar{v}),H(\bar{v}))$, $d^{\lambda}=0$. Therefore, $J_{r}F_{\epsilon}(\bar{r})$ is nonsingular. ∎
@@
Let $r^{k}$} be generated by Algorithm 1. Suppose Assumption 5.5 holds at the accumulation point $\bar{r}$ of $r^{k}$}, then (a) $\nabla g_{\epsilon}(\bar{r})=0$. (b) If the accumulation point $\bar{r}$ satisfies $F_{\epsilon}(\bar{r})=0$, then $r^{k}$} converge to $\bar{r}$ Q-quadratically.
@@
To close this section, we designed the damped Newton method to solve the subproblem (NLPϵ) by its KKT conditions (20). Theorem 3.1 guarantees SDNM converges to an C-stationary point under MPEC-LICQ, and Theorem 5.8 implies that the damped Newton method enjoys a quadratic convergence rate under Assumption 5.5.
@@ 6 Numerical Experiments
1:Given $T$, split the data set into a subset $\Omega$ with $p_{1}$ points and a hold-out test set $\Theta$ with $p_{2}$ points. The set $\Omega$ is equally partitioned into $T$ pairwise disjoint subsets, one for each fold.
@@
3:Post-processing procedure. The regularization hyperparameter $\hat{C}$ is rescaled by a factor T/(T1)T/(T-1). This gives the final classifier $\hat{w}$.
@@ 6.1 The performance of SDNM
In this subsection, we discuss about SDNM from the following two aspects: (i) The decrease of $\|F_{\epsilon}(r^{k})\|_{2}$; (ii) Verification of Assumption 5.5.
6.1.1 The decrease of $\|F_{\epsilon}(r^{k})\|_{2}$
Figure 1 demonstrates the typical decrease of $\|F_{\epsilon}(r^{k})\|_{2}$ for some data sets, where one can indeed observe the quadratic convergence rate of SDNM. Each $\star$ indicates the end of an inner loop.
Figure 1: $\log_{10}\|F_{\epsilon}(r^{k})\|_{2}$ along iterations by SDNM
@@
One can observe from Table 2 that for most data set, the outer loop reaches the stopping criteria within 30 iterations by SDNM. We define $A_{2}:=\bar{r}^{\top}_{v}\nabla^{2}_{vv}L_{\epsilon}(\bar{v},\bar{\lambda})\bar{r}_{v}$, and it can be seen that $A_{2}$ is always positive, which satisfies Assumption 5.5, implying that for each $\epsilon$, the damped Newton method provides strict local minimizers of (NLPϵ).
@@
No. Data set CC\ $A_{2}\$$(k,iter)$ No. Data set CC\ $A_{2}\$$(k,iter)$
1 diabetes 2.93 1483.76 $(\mathbf{13},1064)$ 7 phishing 1.77 31.81 $(\mathbf{13},730)$
2 heart 0.50 6.49 $(\mathbf{10},1123)$ 8 w1a 0.03 43.86 $(\mathbf{10},458)$
3 australian 3.13 30.65 $(\mathbf{10},1248)$ 9 w2a 5.20 488.09 $(\mathbf{11},582)$
4 german.number 15.00 997.64 $(\mathbf{25},619)$ 10 w3a 0.08 87.95 $(\mathbf{18},2119)$
5 ionosphere 8.55 889.06 $(\mathbf{12},915)$ 11 w4a 2.31 4761.02 $(\mathbf{14},1391)$
6 sonar 7.82 428.05 $(\mathbf{15},939)$ 12 w5a 0.07 58.98 $(\mathbf{20},1871)$
6.2 Comparison with other methods
We set the initial values of the parameters of SDNM as $C=1,\ \rho=0.5,\ \sigma=1\times{10}^{-4}$. We compare our method with the following two methods:
  • fmincon, solve (MPEC) directly with a Matlab built-in function fmincon.
  • SGRM, solve (MPEC) by Scholtes global relaxation method proposed in [23]. SNOPT [15] is used to solve the subproblem in each iteration $k$ (given $s_{k}>0$):
@@ 6.2 Comparison with other methods
    CPU time, which is the time of hyperparameter optimization solving by aforementioned methods with the given data set in Table 1.
@@
    Test error ($E_{te}\left(\%\right)$) as defined by $E_{te}=\frac{1}{p_{2}}\sum_{(x_{i},y_{i})\in\Theta}\lvert{\rm sign}\left(x_{i}^{\top}w^{*}\right)-y_{i}\rvert$, where $\Theta$ is the test set.
@@
The results are reported in Table 3, where we mark the winners of $t$, $E_{te}\left(\%\right)$ and $E_{CV}\left(\%\right)$ in bold. ($\beta_{1}$$\beta_{2}$) represents the number of dimensions of variable and the number of constraints in the optimization problem of each methods. As shown in Table 3, SDNM performs better than other methods in terms of test error ($E_{te}(\%)$) and CV error ($E_{CV}\left(\%\right)$), implying that this method has superior generalization performance on classification prediction and solving the over-fitting problem.
In terms of the CPU time summarized in Figure 2, SDNM takes the shortest time among the three methods in all data sets. In order to demonstrate how much the SDNM accelerates the optimization process, in Table 3, we also present the time ratio, which shows that SNDM is always faster than the better one of the existing approaches (fmincon and SGRM), and is about twice to 20 times as fast as the slower one. In particular, SGRM does well in small-scale data sets (No.1-7), and fmincon is competitive in dealing with large scale data sets (No.8-12), but SDNM proposed in this paper is the winner of all of the datasets. For example, dataset w1a (No.8) with 450 cross-validation data points and 300 features can return a solution of hyperparameter in 10 seconds, which is over 20 times faster than SGRM and over 3 times faster than fmincon.
@@
No. Dataset Method $(\beta_{1},\beta_{2})$ $C$ CPU time time ratio $E_{te}$ $E_{CV}$
1 diabetes fmincon $(1801,5400)$ 9.80 16.5 1.99 $\mathbf{23.5}$ $\mathbf{22.7}$
SGRM $(1801,5400)$ 1.50 8.7 1.05 23.7 23.7
SDNM $(1801,1800)$ 2.93 $\mathbf{8.3}$ 1.00 24.1 23.0
2 heart fmincon $(901,2700)$ 11.81 3.4 2.90 18.3 $\mathbf{16.0}$
SGRM $(901,2700)$ 1.62 1.7 1.45 17.5 18.0
SDNM $(901,900)$ 0.50 $\mathbf{1.2}$ 1.00 $\mathbf{15.8}$ 18.0
3 australian fmincon $(1801,5400)$ 4.83 17.7 2.32 $\mathbf{14.6}$ $\mathbf{13.3}$
SGRM $(1801,5400)$ 1.46 15.9 2.09 $\mathbf{14.6}$ 14.0
SDNM $(1801,1800)$ 3.13 $\mathbf{7.6}$ 1.00 15.1 13.7
4 german. fmincon $(1801,5400)$ 25.15 18.3 2.08 25.3 $\mathbf{26.7}$
number SGRM $(1801,5400)$ 1.40 12.1 1.38 $\mathbf{24.7}$ 28.7
SDNM $(1801,1800)$ 15.00 $\mathbf{8.8}$ 1.00 25.3 $\mathbf{26.7}$
5 ionosphere fmincon $(1441,4320)$ 8.76 11.6 2.23 $\mathbf{2.7}$ 26.7
SGRM $(1441,4320)$ 1.50 5.5 1.06 4.5 $\mathbf{25.4}$
SDNM $(1441,1440)$ 8.55 $\mathbf{5.2}$ 1.00 $\mathbf{2.7}$ 26.7
6 sonar fmincon $(901,2700)$ 13.39 4.4 2.82 $\mathbf{27.6}$ 30.0
SGRM $(901,2700)$ 2.64 2.8 1.77 $\mathbf{27.6}$ 30.0
SDNM $(901,900)$ 7.82 $\mathbf{1.6}$ 1.00 $\mathbf{27.6}$ $\mathbf{29.3}$
7 phishing fmincon $(1801,5400)$ 5.64 18.0 3.08 $\mathbf{6.4}$ $\mathbf{7.0}$
SGRM $(1801,5400)$ 1.54 17.4 2.97 $\mathbf{6.4}$ 8.3
SDNM $(1801,1800)$ 1.77 $\mathbf{5.9}$ 1.00 6.9 9.0
8 w1a fmincon $(2701,8100)$ 3.46 35.7 3.93 3.5 $\mathbf{3.6}$
SGRM $(2701,8100)$ 1.50 231.6 25.52 3.3 3.8
SDNM $(2701,2700)$ 0.03 $\mathbf{9.1}$ 1.00 $\mathbf{3.1}$ 3.8
9 w2a fmincon $(2701,8100)$ 3.46 36.1 1.91 $\mathbf{3.2}$ 3.6
SGRM $(2701,8100)$ 1.51 137.1 7.26 $\mathbf{3.2}$ 3.6
SDNM $(2701,2700)$ 5.20 $\mathbf{18.9}$ 1.00 3.3 $\mathbf{3.3}$
10 w3a fmincon $(2701,8100)$ 4.06 37.1 1.01 2.4 3.3
SGRM $(2701,8100)$ 1.50 250.8 6.82 2.4 3.3
SDNM $(2701,2700)$ 0.08 $\mathbf{36.8}$ 1.00 $\mathbf{2.1}$ $\mathbf{2.9}$
11 w4a fmincon $(2701,8100)$ 3.64 36.6 1.23 $\mathbf{3.1}$ $\mathbf{3.1}$
SGRM $(2701,8100)$ 1.48 236.0 7.92 $\mathbf{3.1}$ $\mathbf{3.1}$
SDNM $(2701,2700)$ 2.31 $\mathbf{29.8}$ 1.00 $\mathbf{3.1}$ $\mathbf{3.1}$
12 w5a fmincon $(2701,8100)$ 3.66 37.0 1.01 4.1 4.2
SGRM $(2701,8100)$ 1.52 257.2 7.02 4.1 4.4
SDNM $(2701,2700)$ 0.07 $\mathbf{36.6}$ 1.00 $\mathbf{2.4}$ $\mathbf{3.1}$
Figure 2: CPU time of three methods